# Partial Regularity for the three-dimensional Stochastic Ericksen–Leslie equations

### v1

@@ 4.2. Local-energy-preserving approximation
$\displaystyle\int_{\mathbb{T}^{3}}|\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma})|^{2}{\,\mathrm{d}x}$ $\displaystyle=\int_{\mathbb{T}^{3}}[|\Delta{{\bf{d}}}^{\sigma}|^{2}+|f({{\bf{d}}}^{\sigma})|^{2}-2\Delta{{\bf{d}}}^{\sigma}\cdot f({{\bf{d}}}^{\sigma})]{\,\mathrm{d}x}$
$\displaystyle=\int_{\mathbb{T}^{3}}[|\Delta{{\bf{d}}}^{\sigma}|^{2}+|f({{\bf{d}}}^{\sigma})|^{2}+2\nabla{{\bf{d}}}^{\sigma}:\nabla f({{\bf{d}}}^{\sigma})]{\,\mathrm{d}x}$
$\displaystyle=\int_{\mathbb{T}^{3}}[|\Delta{{\bf{d}}}^{\sigma}|^{2}+|f({{\bf{d}}}^{\sigma})|^{2}-2|\nabla{{\bf{d}}}^{\sigma}|^{2}+2|\nabla{{\bf{d}}}^{\sigma}|^{2}|{{\bf{d}}}^{\sigma}|^{2}+4|(\nabla{{\bf{d}}}^{\sigma})^{\top}{{\bf{d}}}^{\sigma}|^{2}]{\,\mathrm{d}x}.$
By the Gronwall inequality, we can get the uniform energy estimate as follows
(4.11) $\sup_{0<\sigma<1}\sup_{0<t<T}$ $\left(\|{{\bf{v}}}^{\sigma}(t)\|_{H}^{2}+\|{{\bf{d}}}^{\sigma}(t)\|_{H^{1}(\mathbb{T}^{3})}^{2}\right)+\int_{0}^{T}\|\nabla{{\bf{v}}}^{\sigma}(t)\|_{L^{2}(\mathbb{T}^{3})}^{2}+\|\nabla^{2}{{\bf{d}}}^{\sigma}(t)\|_{L^{2}(\mathbb{T}^{3})}^{2}{\,\mathrm{d}t}$
$\leq$ $C\left(\|{\bf{u}}_{0}\|_{H}^{2}+\|{\bf{d}}_{0}\|_{H^{1}(\mathbb{T}^{3})}^{2}+\int_{0}^{T}\|{{\bf{z}}}(t)\|_{L^{4}(\mathbb{T}^{4})}^{4}\,\mathrm{d}t\right)e^{C\int_{0}^{T}\|{\bf{z}}(t)\|_{L^{4}(\mathbb{T}^{3})}^{4}\,\mathrm{d}t}$
=:absent:=:= : $\Psi(T).$
Taking the divergence of (4.9)${}_{1}$ we obtain
(4.12) $-\Delta{\pi}^{\sigma}=\mathop{\rm div}\nolimits^{2}[({{\bf{z}}}+\Phi_{\sigma}[{{\bf{v}}}^{\sigma}])\otimes({{\bf{z}}}+{{\bf{v}}}^{\sigma})]+\mathop{\rm div}\nolimits(\nabla\Phi_{\sigma}[{{\bf{d}}}^{\sigma}]\cdot(\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma}))).$
By the standard elliptic theory and the Sobolev inequalities, we can estimate
$\displaystyle\|{\pi}^{\sigma}\|_{L^{\frac{5}{3}}(Q_{T})}$ $\displaystyle\leq C\|({{\bf{z}}}+\Phi_{\sigma}[{{\bf{v}}}^{\sigma}])\otimes({{\bf{z}}}+{{\bf{v}}}^{\sigma})\|_{L^{\frac{5}{3}}(Q_{T})}+\|\nabla(\Phi_{\sigma}[{{\bf{d}}}^{\sigma}])\cdot(\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma}))\|_{L_{t}^{\frac{5}{3}}L_{x}^{\frac{15}{14}}}$
$\displaystyle\leq C\left(\|{{\bf{z}}}\|_{L^{\frac{10}{3}}(Q_{T})}^{2}+\|{{\bf{v}}}^{\sigma}\|^{2}_{L^{\frac{10}{3}}(Q_{T})}+\|\nabla{{\bf{d}}}^{\sigma}\|^{2}_{L_{t}^{10}L_{x}^{\frac{30}{13}}}\|\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma})\|_{L^{2}(Q_{T})}\right).$
This together with (4.11) implies that
(4.13) $\|{\pi}^{\sigma}\|_{L^{\frac{5}{3}}(Q_{T})}\leq C\Psi(T).$
provided $\int_{\mathbb{T}^{3}}{\pi}^{\sigma}=0.$ Furthermore, we can show with (4.9)${}_{1}$ that
(4.14) $\|\partial_{t}{{\bf{v}}}^{\sigma}\|_{L^{2}(0,T;D(A^{-1}))+L^{2}(0,T;H^{-1}(\mathbb{T}^{3}))+L^{\frac{5}{4}}(Q_{T})}$
$\leq C(\|{{\bf{v}}}^{\sigma}\|_{L^{2}(Q_{T})}+\|{\bf{z}}\|_{L^{4}(Q_{T})}^{2}+\|{{\bf{v}}}^{\sigma}\|_{L^{4}(Q_{T})}^{2}+\|\nabla{\bf{d}}^{\sigma}\|_{L^{\frac{10}{3}}(Q_{T})}\|\Delta{\bf{d}}^{\sigma}-f({{\bf{d}}}^{\sigma})\|_{L^{2}(Q_{T})})$
@@
And from (4.9)${}_{3}$ we can obtain
(4.15) $\|\partial_{t}{\bf{d}}^{\sigma}\|_{L^{\frac{5}{3}}(Q_{T})}$
$\leq C\left(\|{{\bf{z}}}\|_{L^{\frac{10}{3}}(Q_{T})}\|\nabla{\bf{d}}^{\sigma}\|_{L^{\frac{10}{3}}(Q_{T})}+\|{{\bf{v}}}^{\sigma}\|_{L^{\frac{10}{3}}(Q_{T})}\|\nabla{\bf{d}}^{\sigma}\|_{L^{\frac{10}{3}}(Q_{T})}+T^{\frac{1}{10}}\|\Delta{\bf{d}}^{\sigma}-f({{\bf{d}}}^{\sigma})\|_{L^{2}(Q_{T})}\right)$
$\leq C(\Psi(T)+T^{\frac{1}{10}}\sqrt{\Psi(T)}).$

### v2

@@ 4.1.1 Pathwise estimates independent of the approximation parameter
$\displaystyle\left|\int_{\mathbb{T}^{3}}{{\bf{z}}}\cdot\nabla{\Phi_{\sigma}[{{\bf{d}}}^{\sigma}]}\cdot(\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma})){\,\mathrm{d}x}\right|$
$\displaystyle\leq$ $\displaystyle\frac{1}{8}\|\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma})\|_{L^{2}(\mathbb{T}^{3})}^{2}+C\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{2}\|\nabla{{\bf{d}}}^{\sigma}\|_{L^{4}(\mathbb{T}^{3})}^{2},$
$\displaystyle\leq$ $\displaystyle\frac{1}{8}\|\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma})\|_{L^{2}(\mathbb{T}^{3})}^{2}+C\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{2}\|\nabla^{2}{{\bf{d}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{\frac{3}{2}}\|{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{\frac{1}{2}}$
$\displaystyle\leq$ $\displaystyle\frac{1}{8}\|\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma})\|_{L^{2}(\mathbb{T}^{3})}^{2}+\frac{1}{4}\|\nabla^{2}{{\bf{d}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}+C\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{4}\|{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}.$
Combining these two estimations, together with (86), we obtain
$\frac{\,\mathrm{d}}{{\,\mathrm{d}t}}\int_{\mathbb{T}^{3}}\frac{1}{2}(|{{\bf{v}}}^{\sigma}|^{2}+|\nabla{{\bf{d}}}^{\sigma}|^{2})+F({{\bf{d}}}^{\sigma}){\,\mathrm{d}}x+\dfrac{1}{2}\int_{\mathbb{T}^{3}}|\nabla{{\bf{v}}}^{\sigma}|^{2}{\,\mathrm{d}}x$ (22)
$+\dfrac{7}{8}\int_{\mathbb{T}^{3}}|\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma})|^{2}{\,\mathrm{d}}x$
$\leq$ $C\left(\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{4}+{\dfrac{3}{2}}\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{4})}^{4}\|{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}+{\dfrac{1}{2}\|{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}}\right)$
$+\frac{1}{4}\|\nabla^{2}{{\bf{d}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}+C\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{4}\|{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}$
$\leq$ $C\left(\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{4}+{\dfrac{1}{2}\|{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}}\right)+\frac{1}{4}\|\nabla^{2}{{\bf{d}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}+C\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{4}\|{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}.$
Here we note again that the constant C may be different at each occurrence. Now we deal with the last term on the left-hand side (LHS) of (22) as follows
$\displaystyle\int_{\mathbb{T}^{3}}|\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma})|^{2}{\,\mathrm{d}x}$
$\displaystyle=$ $\displaystyle\int_{\mathbb{T}^{3}}[|\Delta{{\bf{d}}}^{\sigma}|^{2}+|f({{\bf{d}}}^{\sigma})|^{2}-2\Delta{{\bf{d}}}^{\sigma}\cdot f({{\bf{d}}}^{\sigma})]{\,\mathrm{d}x}$
$\displaystyle=$ $\displaystyle\int_{\mathbb{T}^{3}}[|\Delta{{\bf{d}}}^{\sigma}|^{2}+|f({{\bf{d}}}^{\sigma})|^{2}+2\nabla{{\bf{d}}}^{\sigma}:\nabla f({{\bf{d}}}^{\sigma})]{\,\mathrm{d}x}$
$\displaystyle=$ $\displaystyle\int_{\mathbb{T}^{3}}[|\Delta{{\bf{d}}}^{\sigma}|^{2}+|f({{\bf{d}}}^{\sigma})|^{2}-2|\nabla{{\bf{d}}}^{\sigma}|^{2}+2|\nabla{{\bf{d}}}^{\sigma}|^{2}|{{\bf{d}}}^{\sigma}|^{2}+4|(\nabla{{\bf{d}}}^{\sigma})^{\top}{{\bf{d}}}^{\sigma}|^{2}]{\,\mathrm{d}x}.$
By slightly rearranging the terms on the RHS, we arrive at
$\int_{\mathbb{T}^{3}}|\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma})|^{2}{\,\mathrm{d}x}$ (23)
$=$ $\int_{\mathbb{T}^{3}}[|\Delta{{\bf{d}}}^{\sigma}|^{2}+|f({{\bf{d}}}^{\sigma})|^{2}+2|\nabla{{\bf{d}}}^{\sigma}|^{2}|{{\bf{d}}}^{\sigma}|^{2}+4|(\nabla{{\bf{d}}}^{\sigma})^{\top}{{\bf{d}}}^{\sigma}|^{2}]{\,\mathrm{d}x}-2\int_{\mathbb{T}^{3}}|\nabla{{\bf{d}}}^{\sigma}|^{2}{\,\mathrm{d}x}.$
Combining (23) and (22), we obtain, after some slight rearrangements:
$\frac{\,\mathrm{d}}{{\,\mathrm{d}t}}\int_{\mathbb{T}^{3}}\frac{1}{2}(|{{\bf{v}}}^{\sigma}|^{2}+|\nabla{{\bf{d}}}^{\sigma}|^{2})+F({{\bf{d}}}^{\sigma}){\,\mathrm{d}}x+|\nabla{{\bf{v}}}^{\sigma}|^{2}_{L^{2}(\mathbb{T}^{3})}+\dfrac{1}{8}\int_{\mathbb{T}^{3}}|\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma})|^{2}{\,\mathrm{d}x}$ (24)
$+\dfrac{6}{8}\int_{\mathbb{T}^{3}}[|\Delta{{\bf{d}}}^{\sigma}|^{2}+|f({{\bf{d}}}^{\sigma})|^{2}+2|\nabla{{\bf{d}}}^{\sigma}|^{2}|{{\bf{d}}}^{\sigma}|^{2}+4|(\nabla{{\bf{d}}}^{\sigma})^{\top}{{\bf{d}}}^{\sigma}|^{2}]{\,\mathrm{d}x}$
$\leq C\left(\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{4}+{\dfrac{1}{2}\|{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}}\right)+\frac{1}{4}\|\nabla^{2}{{\bf{d}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}+C\|{{\bf{z}}}\|_{L^{4}(\mathbb{T}^{3})}^{4}\|{{\bf{v}}}^{\sigma}\|_{L^{2}(\mathbb{T}^{3})}^{2}$
$+\dfrac{6}{4}\int_{\mathbb{T}^{3}}|\nabla{{\bf{d}}}^{\sigma}|^{2}{\,\mathrm{d}x}.$
By employing techniques including an application of the Gronwall inequality to (24) (for a detailed derivation, we refer the reader to Section C of the Appendix), we obtain
$\sup_{0<t<T}\left(\|{{\bf{v}}}^{\sigma}(t)\|_{H}^{2}+\|{{\bf{d}}}^{\sigma}(t)\|_{H^{1}(\mathbb{T}^{3})}^{2}\right)$ (25)
$+\int_{0}^{T}\|\nabla{{\bf{v}}}^{\sigma}(t)\|_{L^{2}(\mathbb{T}^{3})}^{2}+\|\nabla^{2}{{\bf{d}}}^{\sigma}(t)\|_{L^{2}(\mathbb{T}^{3})}^{2}{\,\mathrm{d}t}$
$\leq$ $C\left(\|{\bf{u}}_{0}\|_{H}^{2}+\|{\bf{d}}_{0}\|_{H^{1}(\mathbb{T}^{3})}^{2}+\int_{0}^{T}\|{{\bf{z}}}(t)\|_{L^{4}(\mathbb{T}^{4})}^{4}\,\mathrm{d}t\right)e^{C\int_{0}^{T}{\frac{1}{2}}+\|{\bf{z}}(t)\|_{L^{4}(\mathbb{T}^{3})}^{4}\,\mathrm{d}t}$
$=$ $C\left(\|{\bf{u}}_{0}\|_{H}^{2}+\|{\bf{d}}_{0}\|_{H^{1}(\mathbb{T}^{3})}^{2}+\|{{\bf{z}}}(t)\|_{L^{4}(Q_{T})}^{4}\right)e^{C\int_{0}^{T}{\frac{1}{2}}+\|{\bf{z}}(t)\|_{L^{4}(\mathbb{T}^{3})}^{4}\,\mathrm{d}t}$
$=:$ $\Psi(T).$
Then from (25) we infer (21), and also we obtain
$\sup_{0<t<T}\left(\|{{\bf{v}}}^{\sigma}(t)\|_{H}^{2}+\|{{\bf{d}}}^{\sigma}(t)\|_{H^{1}(\mathbb{T}^{3})}^{2}\right)+\|\nabla{{\bf{v}}}^{\sigma}\|_{L^{2}(Q_{T})}^{2}+\|\nabla^{2}{{\bf{d}}}^{\sigma}\|_{L^{2}(Q_{T})}^{2}\leq\Psi(T),$
which implies (20)1. Also from (24) and (25) we obtain (20)5.
Next, we aim to derive (20)2, a uniform bound for ${\pi}^{\sigma}$ in $\sigma$. Taking the divergence of (17)1 we obtain
$\Delta{\pi}^{\sigma}=\mathop{\rm div}\nolimits^{2}[({{\bf{z}}}+\Phi_{\sigma}[{{\bf{v}}}^{\sigma}])\otimes({{\bf{z}}}+{{\bf{v}}}^{\sigma})]+\mathop{\rm div}\nolimits(\nabla\Phi_{\sigma}[{{\bf{d}}}^{\sigma}]\cdot(\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma}))).$ (26)
Applying the elliptic theory as in Lemma 5.1 of [78], and utilizing Hölder’s inequality we obtain:
$\|{\pi}^{\sigma}\|_{L^{\frac{5}{3}}(Q_{T})}\leq$ $C\|({{\bf{z}}}+\Phi_{\sigma}[{{\bf{v}}}^{\sigma}])\otimes({{\bf{z}}}+{{\bf{v}}}^{\sigma})\|_{L^{\frac{5}{3}}(Q_{T})}$ (27)
$+\|\nabla(\Phi_{\sigma}[{{\bf{d}}}^{\sigma}])\cdot(\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma}))\|_{L^{\frac{5}{3}}(0,T;W^{-1,\frac{5}{3}}(\mathbb{T}^{3}))}$
[…]
$+\|\nabla(\Phi_{\sigma}[{{\bf{d}}}^{\sigma}])\cdot(\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma}))\|_{L^{\frac{5}{3}}(0,T;W^{-1,\frac{5}{3}}(\mathbb{T}^{3}))}.$
Now we first deal with the first term on the RHS. We easily observe that
$\|{{\bf{z}}}\|_{L^{\frac{10}{3}}(Q_{T})}\leq\|{{\bf{z}}}\|_{L^{\infty}(Q_{T})}.$ (28)
By the Riesz–Thorin theorem and the Sobolev embedding theorem in 3D we observe that
$\|{{\bf{v}}}^{\sigma}\|_{L^{\frac{10}{3}}(Q_{T})}$ $\leq C\|{{\bf{v}}}^{\sigma}\|_{L^{2}(0,T;L^{6}(\mathbb{T}^{3}))}^{\frac{2}{5}}\|{{\bf{v}}}^{\sigma}\|_{L^{\infty}(0,T;L^{2}(\mathbb{T}^{3}))}^{\frac{3}{5}}$ (29)
$\leq C\|{{\bf{v}}}^{\sigma}\|_{L^{2}(0,T;H^{1}(\mathbb{T}^{3}))}^{\frac{2}{5}}\|{{\bf{v}}}^{\sigma}\|_{L^{\infty}(0,T;L^{2}(\mathbb{T}^{3}))}^{\frac{3}{5}}$
Then this together with (28) and (27) imply that
$\|{\pi}^{\sigma}\|_{L^{\frac{5}{3}}(Q_{T})}\leq$ $C\left(\|{{\bf{z}}}\|_{L^{\infty}(Q_{T})}^{2}+C\|{{\bf{v}}}^{\sigma}\|_{L^{2}(0,T;H^{1}(\mathbb{T}^{3}))}^{\frac{4}{5}}\|{{\bf{v}}}^{\sigma}\|_{L^{\infty}(0,T;L^{2}(\mathbb{T}^{3}))}^{\frac{6}{5}}\right)$ (30)
$+\|\nabla(\Phi_{\sigma}[{{\bf{d}}}^{\sigma}])\cdot(\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma}))\|_{L^{\frac{5}{3}}(0,T;W^{-1,\frac{5}{3}}(\mathbb{T}^{3}))}.$
Next we deal with the second term on the RHS of (30). We first observe that in 3D, the space $L^{\frac{15}{14}}(\mathbb{T}^{3})$ is embedded in $W^{-1,\frac{5}{3}}(\mathbb{T}^{3})$ thanks to the Sobolev inequalities , so applying this observation to the second term on the RHS we obtain
$\|{\pi}^{\sigma}\|_{L^{\frac{5}{3}}(Q_{T})}\leq$ $C\left(\|{{\bf{z}}}\|_{L^{\infty}(Q_{T})}^{2}+C\|{{\bf{v}}}^{\sigma}\|_{L^{2}(0,T;H^{1}(\mathbb{T}^{3}))}^{\frac{4}{5}}\|{{\bf{v}}}^{\sigma}\|_{L^{\infty}(0,T;L^{2}(\mathbb{T}^{3}))}^{\frac{6}{5}}\right)$ (31)
$+\|\nabla(\Phi_{\sigma}[{{\bf{d}}}^{\sigma}])\cdot(\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma}))\|_{L^{\frac{5}{3}}(0,T;L^{\frac{15}{14}}(\mathbb{T}^{3}))}.$
Now applying Hölder’s inequality to the second term on the RHS we obtain
$\|{\pi}^{\sigma}\|_{L^{\frac{5}{3}}(Q_{T})}\leq$ $C\left(\|{{\bf{z}}}\|_{L^{\infty}(Q_{T})}^{2}+C\|{{\bf{v}}}^{\sigma}\|_{L^{2}(0,T;H^{1}(\mathbb{T}^{3}))}^{\frac{4}{5}}\|{{\bf{v}}}^{\sigma}\|_{L^{\infty}(0,T;L^{2}(\mathbb{T}^{3}))}^{\frac{6}{5}}\right)$ (32)
$+C\|\nabla{{\bf{d}}}^{\sigma}\|^{2}_{L^{10}(0,T;L^{\frac{30}{13}}(\mathbb{T}^{3})}\|\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma})\|_{L^{2}(Q_{T})}.$
Now applying Riesz–Thorin theorem to the second term on the RHS of (32), and utilizing the Sobolev embedding theorem in 3D, we observe that
$\|\nabla{{\bf{d}}}^{\sigma}\|_{L^{10}(0,T;L^{\frac{30}{13}}(\mathbb{T}^{3}))}\leq C\|\nabla{{\bf{d}}}^{\sigma}\|_{L^{\infty}(0,T;L^{2}(\mathbb{T}^{3}))}^{\frac{1}{5}}\times\|\nabla{{\bf{d}}}^{\sigma}\|_{L^{2}(0,T;L^{6}(\mathbb{T}^{3}))}^{\frac{4}{5}}$ (33)
$\leq C\|\nabla{{\bf{d}}}^{\sigma}\|_{L^{\infty}(0,T;L^{2}(\mathbb{T}^{3}))}^{\frac{1}{5}}\times\|\nabla{{\bf{d}}}^{\sigma}\|_{L^{2}(0,T;H^{1}(\mathbb{T}^{3}))}^{\frac{4}{5}}$
This together with (32) implies that
$\|{\pi}^{\sigma}\|_{L^{\frac{5}{3}}(Q_{T})}$ (34)
$\leq$ $C\left(\|{{\bf{z}}}\|_{L^{\infty}(Q_{T})}^{2}+C\|{{\bf{v}}}^{\sigma}\|_{L^{2}(0,T;H^{1}(\mathbb{T}^{3}))}^{\frac{4}{5}}\|{{\bf{v}}}^{\sigma}\|_{L^{\infty}(0,T;L^{2}(\mathbb{T}^{3}))}^{\frac{6}{5}}\right)$
$+C\|\nabla{{\bf{d}}}^{\sigma}\|_{L^{\infty}(0,T;L^{2}(\mathbb{T}^{3}))}^{\frac{2}{5}}\|\nabla{{\bf{d}}}^{\sigma}\|_{L^{2}(0,T;H^{1}(\mathbb{T}^{3}))}^{\frac{8}{5}}\|\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma})\|_{L^{2}(Q_{T})}$
Finally from (34), (20)1, and (20)5 we obtain
$\|{\pi}^{\sigma}\|_{L^{\frac{5}{3}}(Q_{T})}\leq C\Psi(T)+{C\|{{\bf{z}}}\|_{L^{\infty}(Q_{T})}^{2}},$ (35)
which is (20)4.
Next, to prove (20)3, we can utilize (17)1 to show that
$\|\partial_{t}{{\bf{v}}}^{\sigma}\|_{L^{2}(0,T;D(A^{-1}))+L^{2}(0,T;H^{-1}(\mathbb{T}^{3}))+L^{\frac{5}{4}}(Q_{T})}$ (36)
$\leq$ $C(\|{{\bf{v}}}^{\sigma}\|_{L^{2}(Q_{T})}+\|{\bf{z}}\|_{L^{4}(Q_{T})}^{2}+\|{{\bf{v}}}^{\sigma}\|_{L^{4}(Q_{T})}^{2}$
$+\|\nabla{\bf{d}}^{\sigma}\|_{L^{\frac{10}{3}}(Q_{T})}\|\Delta{\bf{d}}^{\sigma}-f({{\bf{d}}}^{\sigma})\|_{L^{2}(Q_{T})})$
@@
Finally, to prove (20)4, we can utilize (17)3 to derive that
$\|\partial_{t}{\bf{d}}^{\sigma}\|_{L^{\frac{5}{3}}(Q_{T})}$ (37)
$\leq$ $C\|{{\bf{z}}}\|_{L^{\frac{10}{3}}(Q_{T})}\|\nabla{\bf{d}}^{\sigma}\|_{L^{\frac{10}{3}}(Q_{T})}+C\|{{\bf{v}}}^{\sigma}\|_{L^{\frac{10}{3}}(Q_{T})}\|\nabla{\bf{d}}^{\sigma}\|_{L^{\frac{10}{3}}(Q_{T})}$
$+CT^{\frac{1}{10}}\|\Delta{\bf{d}}^{\sigma}-f({{\bf{d}}}^{\sigma})\|_{L^{2}(Q_{T})}$
$\leq$ $C(\Psi(T)+CT^{\frac{1}{10}}\sqrt{\Psi(T)}).$
4.2 Approaching the existence of martingale suitable solutions
Based on the pathwise estimates derived above, we can infer a compactness argument from tightness properties of the approximation sequence. Next we will use the Skorokhod embedding theorem (see Theorem 2.4 in [16]) to construct an approximation sequence with a possible shift of the stochastic basis. It is then necessary to verify that the key energy equality (81) is preserved by the constructed sequence. Finally passing to the limit in (17) we obtain the global existence of martingale suitable weak solutions satisfying the local energy inequalities as desired.
4.2.1 Compactness of the approximation sequence